OpenStudy (tabithax3):

I need help using Pascal's triangle to expand the expression: (1/x - √x)^5

OpenStudy (tabithax3):

this is the equation: \[(\frac{ 1 }{ x } - \sqrt{x})^{5}\]

OpenStudy (tabithax3):

@jdoe0001 i got that. i'm confused on how to calculate the square roots after expanding the equation

OpenStudy (campbell_st):

do you know about combinations..?

OpenStudy (tabithax3):

@campbell_st i know \[(a+b)^{5} = a ^{5}+5a ^{4}b+10a ^{3}b ^{2}+10a ^{2}b ^{3}+5ab ^{4}+b ^{5}\]

OpenStudy (campbell_st):

because combinations are used to get to coefficients of each term \[(a + b)^5 = ^5C_{0} (a)^{5 - 0}(b)^0 + ^5C_{1}a^{5 -1}{b}^1 +.....\] in general its \[(a + b)^n = \sum {^nC_{r} (a)^{n - r}b^r}\] where r = 0, 1, 2, 3, 4, ... n

OpenStudy (campbell_st):

that way you don't need to worry about pascals triangle and you should write your terms in index form... easier to calculate \[(x^{-1} - x^{\frac{1}{2}})^5 = ^5C_{0}(x^{-1})^5(x^{\frac{1}{2}})^0 + ^5C_{1} (x^{-1})^{5 -1}(x^{\frac{1}{2}})^1+...\]

OpenStudy (jdoe0001):

so, what part is the one you need help with?

OpenStudy (jdoe0001):

it's not the coefficients, I can see that

OpenStudy (tabithax3):

@jdoe0001 i got \[\frac{ 1 }{ x ^{5} }+\frac{ 5 }{ x ^{4} }(-\sqrt{x})+\frac{ 10 }{x ^{3} }(-\sqrt{x})^{2}+\frac{ 10 }{ x ^{2} }(-\sqrt{x})^{3}+\frac{ 5 }{ x }(-\sqrt{x})^{4}+(-\sqrt{x})^{5}\]

OpenStudy (tabithax3):

@jdoe0001 i got lost right there

OpenStudy (jdoe0001):

ok

OpenStudy (campbell_st):

keep it in index form until you have a final term \[x^-4 \times x^{\frac{1}{2}} = x^{-\frac{7}{2}} = \frac{1}{\sqrt{x^7}}\]

OpenStudy (tabithax3):

the answer is \[\frac{ 1 }{ x ^{5} }-\frac{ 5 }{ x ^{\frac{ 7 }{ 2 }} }+\frac{ 10 }{ x ^{2} }+\frac{ 10 }{ x ^{\frac{ 1 }{ 2 }} }+5x-x ^{\frac{ 5 }{ 2 }}\] but i don't know how to get it from there

OpenStudy (campbell_st):

nope... because you would need to use your index laws skills to write the terms correctly...

OpenStudy (jdoe0001):

$$ \cfrac{1}{x^5}+\cfrac{5\times -\sqrt{x}}{x^5} \implies \cfrac{1+(x(-5\sqrt{x}))}{x^5}\\ \boxed{\cfrac{1-5x\sqrt{x}}{x^5}} $$

OpenStudy (campbell_st):

and I think you have the sign of the 4th term incorrect...

OpenStudy (jdoe0001):

which is just plain fraction addition :|

OpenStudy (tabithax3):

i was just writing what my book wrote. i know the fourth term should be negative/minus

OpenStudy (jdoe0001):

$$ \cfrac{1}{x^5}+\cfrac{5\times -\sqrt{x}}{x^4} \implies \cfrac{1+(x(-5\sqrt{x}))}{x^5}\\ \boxed{\cfrac{1-5x\sqrt{x}}{x^5}} $$ had a typo on an exponent :(

OpenStudy (campbell_st):

well if the book allows factional indices... then thats what you write

OpenStudy (tabithax3):

@campbell_st i never learned the index form in class. how do you do it?

OpenStudy (campbell_st):

wow... its really hard to do binomial theorem work without index law skills... its a fundamental here is a link with some explanations. Its basically a Year 8 piece of work. http://www.staff.vu.edu.au/mcaonline/units/indices/indexlaws.html http://mathematics.laerd.com/maths/indices-intro.php hope this helps

OpenStudy (tabithax3):

@campbell_st thank you

OpenStudy (jdoe0001):

basically, you'd end up just working primarily with the roots and values of x, for say as example in the 5th term: \( (- \sqrt{x^4}) \implies (-1 \times \sqrt{x^4}) \implies -1 \times x^{4/2} \implies -x^2 \)

OpenStudy (tabithax3):

@jdoe0001 i barely got that haha, thank you so much